OpenStudy (anonymous):

Determine how much money will be in a savings account if the initial deposit was $1000 and the interest rate is 3% compounded continuously for 5 years, 10 months. (Round your answer to the nearest cent.)

OpenStudy (anonymous):

3% a year compounded continuously, use \[1000\times e^{.03\times 5\tfrac{5}{6}}\]

OpenStudy (anonymous):

How would I put that into my calculator?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1000 \times+e^{.03\times+5\tfrac{5}{6}}

OpenStudy (anonymous):

copy and paste, link didn't work

OpenStudy (anonymous):

If your calculator has a \[e^{x}\] function, type the number that stands as x and then hit that function. This will raise e to the specified power.